Verify rolle`s theorem for f(x) = sinx +cosx-1 , x € [ 0,90]

Grade:12th pass

1 Answers

akhilesh kondhalkar
6 years ago
So, for verification of Rolles theorem f(a) = f(b) 
here f(a)=f(0)=sin0+cos0-1= 0+1-1= 0
now f(b) =f(90)= sin90+ cos90-1=1+0-1=0
thus for x=[0,90] Rolles theorem is verified